# A solution to the Cauchy dual subnormality problem for a cyclic analytic $2$-isometry with defect operator of rank two

### v1

@@ Abstract.
The Cauchy dual subnormality problem (for short, CDSP) asks whether the Cauchy dual of a $2$-isometry is subnormal. In this article, we prove that if $\mu$ is a linear combination of unit point mass measures at two non-antipodal points on the unit circle, then the Cauchy dual $M_{z}^{\prime}$ of the multiplication operator $M_{z}$ on the Dirichlet space $D(\mu)$ is not subnormal. If the two points are antipodal then the subnormality of the said operator has been already established in the literature. Thus, we have a complete solution of CDSP in this case.
@@
Cauchy dual; Subnormal Operator; de Branges–Rovnyak space; Dirichlet type spaces
@@ 1. Introduction
In this article, we present a solution to the Cauchy dual subnormality problem (for short, CDSP) for cyclic, analytic 2-isometries whose defect operator is of rank two.
@@
An excellent illustration of such a duality is provided by contractive subnormal operators and completely hyperexpansive operators. The interplay between these two classes of operators can be thought of as an operator-theoretic manifestation of two special classes of functions, viz. completely monotone functions and completely alternating functions. This fruitful synthesis of harmonic analysis with operator theory has been extensively explored in the works of Athavale et al. [8, 9, 10, 11, 37], as well as in the works of Jablonski and Stochel [27, 24, 25, 26, 28]. It turns out that the notion of Cauchy dual fits naturally in these circumstances and helps in understanding the subtle interconnections.
The said theme begins with a simple observation :- If $T:\{\alpha_{n}\}$ is a weighted shift operator, then the Cauchy dual $T^{\prime}$ is the weighted shift with weight sequence $\Big\{\displaystyle\frac{1}{\alpha_{n}}\Big\}$. Combining this fact with a result in [14, Chapter 4, Proposition 6.10] enabled Athavale to prove that the Cauchy dual of a completely hyperexpansive weighted shift is a contractive subnormal weighted shift [6]. In the light of this observation, we naturally have the following question:
@@
This question appears in [15, Question 2.11]. Some special cases of this problem have been dealt with in recent years. Chavan et. al. have extensively contributed in solving the problem in some special cases. The phrase ‘Cauchy dual subnormality problem’ (CDSP, for short) has evolved through these considerations.
We now summarize the present status of the problem.
The following results indicate that the evidence is in the positive direction:
@@
    The Cauchy dual of a completely hyperexpansive weighted shift is a subnormal weighted shift. See [6]
@@
    The Cauchy dual of a concave operator is a hyponormal contraction. See [35].
@@
    The Cauchy dual of a $\Delta_{T}$-regular 2-isometry is subnormal. See [12].
@@
    The Cauchy dual of an expansive operator of class Q is a subnormal [17].
@@
    There exist weighted shifts on directed trees whose Cauchy dual is not subnormal. See [4].
@@
Is the Cauchy dual of a 2-isometric operator a subnormal operator?
@@
Here, $P_{\mu}(z)$ is the Poisson integral for measure $\mu$ given by $\int_{\mathbb{T}}\displaystyle\frac{1-|z|^{2}}{|z-\zeta|^{2}}d\mu(\zeta)$,
dAdA denotes the normalized Lebesgue area measure on the open unit disc $\mathbb{D}$. The space $D(\mu)$ is a reproducing kernel Hilbert space and multiplication by the coordinate function $z$, denoted by $M_{z}$, turns out to be a bounded linear operator (see [2, 32]) on $D(\mu)$. For an elaborate discussion on Dirichlet-type spaces, the reader may also consult [22]. S. Richter described a model for a cyclic, analytic $2$-isometry[32]. Indeed, a cyclic, analytic $2$-isometry is unitarily equivalent to a multiplication operator $M_{z}$ on a Dirichlet-type space $D(\mu)$ for some finite positive Borel measure $\mu$ on unit circle $\mathbb{T}$. This model for cyclic analytic 2-isometries and a decomposition theorem of 2-isometries allow us to reformulate Question 1.3 in the following manner:
@@
Classify all finite positive Borel measures on the unit circle for which the Cauchy dual $M_{z}^{\prime}$ of the multiplication operator $M_{z}$ on the Dirichlet type space $D(\mu)$ is subnormal.
@@
Is it true that for all cyclic $2$-isometries $T$ for which $T^{*}T-I$ is of finite rank, the Cauchy dual of $T$ is subnormal?
In fact, in the same article [16], the authors constructed an example of a cyclic, analytic $2$-isometry $T$ for which $T^{*}T-I$ is of rank two whose Cauchy dual is subnormal. The example certainly strengthened the evidence for the affirmative answer to Question 1.5. However, recently an example of a cyclic, analytic $2$-isometry $T$ for which $T^{*}T-I$ is of rank two whose Cauchy dual is not subnormal [30] has been found.
At this stage, the reader is referred to the fact [16, Theorem 6.1] which states that for a cyclic, analytic $2$-isometry $T$ in $\mathcal{B}(H)$, the rank of $T^{*}T-I$ is finite if and only if there exists a finitely supported measure $\mu$ on the unit circle $\mathbb{T}$ such that $T$ is unitarily equivalent to the operator $M_{z}$ on $D(\mu)$. In fact, if the rank of $T^{*}T-I$ is a positive integer $k,$ then the corresponding measure $\mu$ is supported at exactly $k$ points on the unit circle $\mathbb{T}.$ In view of this result, a construction of a cyclic, analytic $2$-isometry defect operator of finite rank reduces to choosing finitely many points on the unit circle and looking at the multiplication operator $M_{z}$ on the Dirichlet space $D(\mu)$, where $\mu$ is supported at the chosen points. Further, in [16, Theorem 6.4] it is claimed that this space $D(\mu)$ coincides with the de Branges-Rovnyak space $H(B)$ for some vector-valued rational function $B$ with simple poles and equality of norms.
@@
  • An example of a cyclic, analytic $2$-isometry $T$ such that $T^{*}T-I$ is of rank two and whose Cauchy dual is subnormal has been constructed in [16, Example 7.2] by choosing a measure on the unit circle supported at $\{1,-1\}$.
@@
In view of the examples cited above, the question 1.4 can be narrowed down to the following problem:
Problem 1 :
Classify all finitely supported positive Borel measures on the unit circle for which the Cauchy dual $M_{z}^{\prime}$ of the multiplication operator $M_{z}$ on the Dirichlet type space $D(\mu)$ is subnormal and $M_{z}^{*}M_{z}-I$ is of rank two.
The present article is devoted to the solution of the above problem. The main theorem proved in this paper is as follows.
@@
Let $\zeta_{1},\zeta_{2}$ be two distinct points on the unit circle. Let $\mu$ be a positive Borel measure of the form $\mu=\delta_{\zeta_{1}}+\delta_{\zeta_{2}}.$ Then the Cauchy dual of $M_{z}$ on $D(\mu)$ is subnormal if and only if the points $\zeta_{1}$ and $\zeta_{2}$ are antipodal.
The proof of the main theorem is presented in section 3. Note that in view of [16, Proposition 7.1], it is sufficient to prove the theorem in the case where the measure $\mu$ is supported on $\{1,\xi\}$ for any point $\xi$ on the unit circle. In the process, we shall obtain the reproducing kernel for the space $D(\mu)$ where $\mu$ is supported on $\{1,\xi\}$.
@@ 2. Preliminaries
    It is a well known fact that $H^{2}$ is a reproducing kernel Hilbert space with the inner product defined as $\langle f,g\rangle:=\displaystyle\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}$, if $f(z)=\displaystyle\sum_{n=0}^{\infty}a_{n}z^{n}$ , $g(z)=\displaystyle\sum_{n=0}^{\infty}b_{n}z^{n}$ and a kernel (called as Szego Kernel) function $K_{\lambda}(z)=\displaystyle\frac{1}{1-\overline{\lambda}z}$. The interested reader may consult [21].
@@
    Dirichlet-type spaces: Note that Dirichlet-type spaces have been described in the previous section. D. Sarason [33] initiated the study of Dirichlet type spaces by way of identifying such a space with a de Branges-Rovnyak space. This association has been further strengthened in recent years in the works of [16, 18, 29]. The identification of a Dirichlet space as a de Branges-Rovnyak space allows one to compute the reproducing kernel for the Dirichlet space.
@@
An operator $T\in\mathcal{B}(H)$ is said to be subnormal if there exist a Hilbert space $\mathcal{K}$ and an operator $S\in\mathcal{B}(K)$ such that $\mathcal{H}\subseteq K$, $S$ is normal and $S|_{\mathcal{H}}=T$. Readers are encouraged to refer [19] for a detailed study of subnormal operators. Agler proved in [1] that $T\in\mathcal{B}(H)$ is a subnormal contraction (means $T$ is subnormal and $\|T\|\leq 1$) if and only if
@@
Following [7], an operator $T\in B(\mathcal{H})$ is said to be completely hyperexpansive if
$B_{n}(T)\leq 0~~\text{for all integers }n\geq 1.$
@@ 2.1. Reproducing kernel of $D(\mu)$
The algorithm for obtaining reproducing kernel of $D(\mu)$ turns out to be a fruitful computational tool for checking the subnormality of the Cauchy dual of $M_{z}$ on $D(\mu)$. We take up the task of computing the reproducing kernel for the space $D(\mu)$ where $\mu$ is the regular, positive, Borel measure on the unit circle $\mathbb{T}$ of the form $\delta_{1}+\delta_{\xi}.$ The technique used in obtaining the reproducing kernel heavily relies on the work of Costara [20]. Thus, we briefly outline the work carried out in [20] and state the results that are used in the sequel. Given $n\in\mathbb{N}$. Let $c_{1},c_{2},\dots,c_{n}$ be positive real numbers and $\zeta_{1},\zeta_{2},\dots,\zeta_{n}$ be distinct points on the unit circle $\mathbb{T}$. Consider the positive Borel measure $\mu=\sum_{j=1}^{n}c_{j}\delta_{\zeta_{j}}$ on $\mathbb{T}$.
As an application of Riesz-Fejer theorem, it has been observed in [20] that there exist $\alpha_{1},\alpha_{2},\dots,\alpha_{n}\in\mathbb{C}\setminus\overline{\mathbb{D}}$ and $d>0$ such that the following equation holds:
$\prod_{j=1}^{n}|z-\zeta_{j}|^{2}+\sum_{j=1}^{n}c_{j}\prod_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}|z-\zeta_{i}|^{2}=d\prod_{j=1}^{n}|z-\alpha_{j}|^{2},~~~z\in\mathbb{T}$ (2.1)
The values of $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ as obtained in equation (2.1) turn out to be important in the expressions for reproducing kernels of two subspaces of $H^{2}$ as obtained in [20, Theorem 3.1, Theorem 4.4]:
@@
    We observe that $O_{\mu}$ is a rational function of degree $n$ having poles outside the closed unit disk and simple zeros at $\zeta_{j}\in\mathbb{T}$, for $j=1,\dots,n$. Then $\prod_{j=1}^{n}(z-\zeta_{j})H^{2}=O_{\mu}H^{2}$ and hence $O_{\mu}H^{2}$ is a closed subspace of $D(\mu)$. The reproducing kernel for $O_{\mu}H^{2}$ is given by (Refer [20, Theorem 3.1])
    $\tilde{K_{\mu}}(z,\lambda)=\displaystyle\frac{O_{\mu}(z)}{1-\overline{\lambda}z}\overline{O_{\mu}(\lambda)}$ (2.3)
@@
    The reproducing kernel for $(O_{\mu}H^{2})^{\perp}$ as a subspace of $D(\mu)$ is given by
    $\hat{K_{\mu}}(z,\lambda)=\displaystyle\sum_{j=1}^{n}g_{j}(\lambda)f_{j}(z)~~~~z,\lambda\in\mathbb{D}$ (2.4)
@@
    $f_{j}(z)=\displaystyle\frac{O_{\mu}(z)}{O_{\mu}^{\prime}(\zeta_{j})(z-\zeta_{j})}$ (2.5)
@@
    $\begin{pmatrix}\overline{g_{1}(\lambda)}\\ \overline{g_{2}(\lambda)}\\ \vdots\\ \overline{g_{n}(\lambda)}\end{pmatrix}=\left(\left[\langle f_{i},f_{j}\rangle_{\mu}\right]_{1\leq i,j\leq n}\right)^{-1}\begin{pmatrix}f_{1}(\lambda)\\ f_{2}(\lambda)\\ \vdots\\ f_{n}(\lambda)\end{pmatrix}$ (2.6)
@@
    The Lemma [20, Lemma 4.2, 4.3] provides formula for obtaining $\langle f_{i},f_{j}\rangle_{\mu}$ for any $1\leq i,j\leq n$ as follows :
    $\|f_{i}\|_{\mu}^{2}=\langle f_{i},f_{i}\rangle_{\mu}=c_{i}\zeta_{i}f_{i}^{\prime}(\zeta_{i})$ (2.7)
@@
    $\langle f_{i},f_{j}\rangle_{\mu}=\displaystyle\frac{1}{O_{\mu}^{\prime}(\zeta_{i})\overline{O_{\mu}^{\prime}(\zeta_{j})}(1-\zeta_{i}\overline{\zeta_{j}})}$ (2.8)
@@
Let $\mu$ be a positive Borel measure on unit circle $\mathbb{T}$ supported on $\{1,\xi\}$ defined as $\mu=\delta_{1}+\delta_{\xi}$. The reproducing kernel for $D(\mu)$ is given by:
@@
$+\displaystyle\frac{(z-\xi)(\bar{\lambda}-1)}{\Delta(\bar{\xi}-1)}+\frac{(z-1)(\bar{\lambda}-\bar{\xi})}{\Delta(\bar{\xi}-1)}\Bigg],$
where $b,\Delta\in\mathbb{R}$ are constants and $q(z)$ is a polynomial of degree $2$ whose roots are in $\mathbb{C}\setminus\overline{\mathbb{D}}$.

### v2

@@ Abstract.
The Cauchy dual subnormality problem (for short, CDSP) asks whether the Cauchy dual of a $2$-isometry is subnormal. In this article, we prove that if $\mu$ is a sum of unit point mass measures at two non-antipodal points on the unit circle, then the Cauchy dual $M_{z}^{\prime}$ of the multiplication operator $M_{z}$ on the Dirichlet-type space $D(\mu)$ is not subnormal. If the points are antipodal then the subnormality of the said operator has been already established in the literature. Thus, we have a complete solution to CDSP in this case.
@@
Cauchy dual; Subnormal Operator; de Branges–Rovnyak space; Dirichlet-type spaces
*Corresponding Author: V. M. sholapurkar. Email: vmshola@gmail.com
@@ 1. Introduction
In this article, we present a solution to the Cauchy dual subnormality problem (for short, CDSP) for a family of cyclic, analytic $2$-isometries $T$ whose defect operator $T^{*}T-I$ is of rank two.
@@
An excellent illustration of such a duality is provided by contractive subnormal operators and completely hyperexpansive operators. The interplay between these two classes of operators can be thought of as an operator-theoretic manifestation of two special classes of functions, viz. completely monotone functions and completely alternating functions. This fruitful synthesis of harmonic analysis with operator theory has been extensively explored in the works of Athavale et al. [10, 6, 7, 8, 35], as well as in the works of Jabłoński and Stochel [26, 23, 24, 22, 25]. It turns out that the notion of Cauchy dual fits naturally in these circumstances and helps in understanding the subtle interconnections.
The said theme begins with an observation: The Cauchy dual of a completely hyperexpansive weighted shift is a contractive subnormal weighted shift [9]. In the light of this observation, we naturally have the following question:
@@
This question appears in [15, Question 2.11]. The question has been extensively studied in the literature. The following results indicate that the evidence is in the positive direction:
@@
    The Cauchy dual of a completely hyperexpansive weighted shift is a subnormal weighted shift [9].
@@
    The Cauchy dual of a concave operator is a hyponormal contraction [34].
@@
    The Cauchy dual of a $\Delta_{T}$-regular 2-isometry is subnormal [11].
@@
    The Cauchy dual of an expansive operator of class Q is subnormal [14].
@@
    There exist weighted shifts on directed trees whose Cauchy dual is not subnormal [4].
@@
Is the Cauchy dual of a cyclic $2$-isometric operator a subnormal operator?
@@
Here, $P_{\mu}(z)$ is the Poisson integral for measure $\mu$ given by $\int_{\mathbb{T}}\displaystyle\frac{1-|z|^{2}}{|z-\zeta|^{2}}d\mu(\zeta)$ and $dA$ denotes the normalized Lebesgue area measure on the open unit disc $\mathbb{D}$. The space $D(\mu)$ is a reproducing kernel Hilbert space and multiplication by the coordinate function $z$, denoted by $M_{z}$, turns out to be a bounded linear operator (see [2, 30]) on $D(\mu)$. For an elaborate discussion on Dirichlet-type spaces, the reader may also consult [20]. S. Richter described a model for a cyclic, analytic $2$-isometry [30]. Indeed, a cyclic, analytic $2$-isometry is unitarily equivalent to a multiplication operator $M_{z}$ on a Dirichlet-type space $D(\mu)$ for some finite positive Borel measure $\mu$ on unit circle $\mathbb{T}$. This model for cyclic analytic 2-isometries and a decomposition theorem of 2-isometries allow us to revisit the the following question that appeared in [13]:
@@
Classify all finite positive Borel measures on the unit circle for which the Cauchy dual $M_{z}^{\prime}$ of the multiplication operator $M_{z}$ on the Dirichlet-type space $D(\mu)$ is subnormal.
@@
Is it true that for all cyclic, analytic $2$-isometries $T$ for which $T^{*}T-I$ is of finite rank, the Cauchy dual of $T$ is subnormal?
In fact, in the same article [13], the authors constructed an example of a cyclic, analytic $2$-isometry $T$ for which $T^{*}T-I$ is of rank two whose Cauchy dual is subnormal. The example certainly strengthened the evidence for the affirmative answer to Question 1.5. However, recently an example of a cyclic, analytic $2$-isometry $T$ for which $T^{*}T-I$ is of rank two whose Cauchy dual is not subnormal has been found in [28].
At this stage, the reader is referred to the fact [13, Theorem 6.1] which states that for a cyclic, analytic $2$-isometry $T$ in $\mathcal{B}(H)$, the rank of $T^{*}T-I$ is finite if and only if there exists a finitely supported measure $\mu$ on the unit circle $\mathbb{T}$ such that $T$ is unitarily equivalent to the operator $M_{z}$ on $D(\mu)$. In fact, if the rank of $T^{*}T-I$ is a positive integer $k,$ then the corresponding measure $\mu$ is supported at exactly $k$ points on the unit circle $\mathbb{T}.$ In view of this result, a construction of a cyclic, analytic $2$-isometry with defect operator of finite rank reduces to choosing finitely many points on the unit circle and looking at the multiplication operator $M_{z}$ on the Dirichlet-type space $D(\mu)$, where $\mu$ is supported at the chosen points. Further, in [13, Theorem 6.4] it is claimed that this space $D(\mu)$ coincides with the de Branges-Rovnyak space $H(B)$ for some vector-valued rational function $B$ with equality of norms.
@@
  • An example of a cyclic, analytic $2$-isometry $T$ such that $T^{*}T-I$ is of rank two and whose Cauchy dual is subnormal has been constructed in [13, Example 7.2] by choosing a measure on the unit circle supported at $\{1,-1\}$.
@@
In view of the examples cited above, the Question 1.4 can be narrowed down to the following question:
Question 1.6.
Classify all finitely supported positive Borel measures on the unit circle for which the Cauchy dual $M_{z}^{\prime}$ of the multiplication operator $M_{z}$ on the Dirichlet-type space $D(\mu)$ is subnormal and $M_{z}^{*}M_{z}-I$ is of rank two.
The present article is devoted to a partial solution of the above question. The main theorem proved in this paper is as follows.
Theorem 1.7 (Main Theorem).
Let $\zeta_{1},\zeta_{2}$ be two distinct points on the unit circle. Let $\mu$ be a positive Borel measure of the form $\mu=\delta_{\zeta_{1}}+\delta_{\zeta_{2}}$ defined on the unit circle $\mathbb{T}$. If $\zeta_{1},\zeta_{2}$ are non-antipodal points, then the Cauchy dual of $M_{z}$ on $D(\mu)$ is not subnormal.
The proof of the main theorem is presented in section 3. Note that in view of [13, Proposition 7.1], it is sufficient to prove the theorem in the case where the measure $\mu$ is supported on $\{1,\xi\}$ for any point $\xi~(\xi\neq-1)$ on the unit circle. In the process, we shall obtain the reproducing kernel for the space $D(\mu)$ where $\mu$ is supported on $\{1,\xi\}$.
Theorem 1.7 coupled with [13, Theorem 2.4] yields the following characterization in this case.
@@
Let $\zeta_{1},\zeta_{2}$ be two distinct points on the unit circle. Let $\mu$ be a positive Borel measure of the form $\mu=\delta_{\zeta_{1}}+\delta_{\zeta_{2}}$ defined on the unit circle $\mathbb{T}$. Then the Cauchy dual of $M_{z}$ on $D(\mu)$ is subnormal if and only if the points $\zeta_{1}$ and $\zeta_{2}$ are antipodal.
@@ 2. Preliminaries
    It is a well known fact that $H^{2}$ is a reproducing kernel Hilbert space with the inner product defined as $\langle f,g\rangle:=\displaystyle\sum_{n=0}^{\infty}a_{n}\overline{b_{n}}$, if $f(z)=\displaystyle\sum_{n=0}^{\infty}a_{n}z^{n}$ , $g(z)=\displaystyle\sum_{n=0}^{\infty}b_{n}z^{n}$ and a kernel (namely Szego Kernel) function $K_{\lambda}(z)=\displaystyle\frac{1}{1-\overline{\lambda}z}$. The interested reader may consult [19].
@@
    Dirichlet-type spaces: Note that Dirichlet-type spaces have been described in the previous section. D. Sarason [31] initiated the study of Dirichlet-type spaces by way of identifying such a space with a de Branges-Rovnyak space. This association has been further strengthened in recent years in the works of [13, 16, 27]. The identification of a Dirichlet-type space as a de Branges-Rovnyak space allows one to compute the reproducing kernel for the Dirichlet-type space.
@@
An operator $T\in\mathcal{B}(H)$ is said to be subnormal if there exist a Hilbert space $\mathcal{K}$ and an operator $S\in\mathcal{B}(K)$ such that $\mathcal{H}\subseteq K$, $S$ is normal and $S|_{\mathcal{H}}=T$. Readers are encouraged to refer [17] for a detailed study of subnormal operators. Agler proved in [1] that $T\in\mathcal{B}(H)$ is a subnormal contraction ($T$ is subnormal and $\|T\|\leq 1$) if and only if
@@
Following [9], an operator $T\in B(\mathcal{H})$ is said to be completely hyperexpansive if $B_{n}(T)\leq 0~~\text{for all integers }n\geq 1.$
@@ 2.1. Reproducing kernel of $D(\mu)$
The algorithm for obtaining reproducing kernel of $D(\mu)$ turns out to be a fruitful computational tool for checking the subnormality of the Cauchy dual of $M_{z}$ on $D(\mu)$. For $\xi\in\mathbb{T}~(\xi\neq\pm 1)$, we take up the task of computing the reproducing kernel for the space $D(\mu)$ where $\mu$ is the regular, positive, Borel measure on the unit circle $\mathbb{T}$ of the form $\delta_{1}+\delta_{\xi},$ where $\delta_{\lambda}$ denotes the unit point mass measure at a point $\lambda\in\mathbb{T}.$ The technique used in obtaining the reproducing kernel heavily relies on the work of Costara [18]. Thus, we briefly outline the work carried out in [18] and state the results that are used in the sequel. Given $n\in\mathbb{N}$, let $c_{1},c_{2},\dots,c_{n}$ be positive real numbers and $\zeta_{1},\zeta_{2},\dots,\zeta_{n}$ be distinct points on the unit circle $\mathbb{T}$. Consider the positive Borel measure $\mu=\sum_{j=1}^{n}c_{j}\delta_{\zeta_{j}}$ on $\mathbb{T}$.
As an application of Riesz-Fejer theorem, it is observed in [18] that there exist $\alpha_{1},\alpha_{2},\dots,\alpha_{n}\in\mathbb{C}\setminus\overline{\mathbb{D}}$ and $d>0$ such that the following equation holds:
(2.1) $\prod_{j=1}^{n}|z-\zeta_{j}|^{2}+\sum_{j=1}^{n}c_{j}\prod_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}|z-\zeta_{i}|^{2}=d\prod_{j=1}^{n}|z-\alpha_{j}|^{2},~~~z\in\mathbb{T}.$
The values of $\alpha_{1},\alpha_{2},\dots,\alpha_{n}$ as obtained in equation (2.1) turn out to be important in the expression for reproducing kernel of $H^{2}.$ Following [18, Theorem 3.1, Theorem 4.4],
@@
    We observe that $O_{\mu}$ is a rational function of degree $n$ having poles outside the closed unit disk and simple zeros at $\zeta_{j}\in\mathbb{T}$, for each $j=1,\dots,n$. Then $\prod_{j=1}^{n}(z-\zeta_{j})H^{2}=O_{\mu}H^{2}$ and hence $O_{\mu}H^{2}$ is a closed subspace of $D(\mu)$. The reproducing kernel for $O_{\mu}H^{2}$ is given by (Refer [18, Theorem 3.1])
    (2.3) $\tilde{K_{\mu}}(z,\lambda)=\displaystyle\frac{O_{\mu}(z)}{1-\overline{\lambda}z}\overline{O_{\mu}(\lambda)}.$
@@
    The reproducing kernel for $(O_{\mu}H^{2})^{\perp}$ as a subspace of $D(\mu)$ is given by (Refer [18, Theorem 4.4])
    (2.4) $\hat{K_{\mu}}(z,\lambda)=\displaystyle\sum_{j=1}^{n}g_{j}(\lambda)f_{j}(z)~~~~z,\lambda\in\mathbb{D}$
@@
    (2.5) $f_{j}(z)=\displaystyle\frac{O_{\mu}(z)}{O_{\mu}^{\prime}(\zeta_{j})(z-\zeta_{j})}$
@@
    (2.6) $\begin{pmatrix}\overline{g_{1}(\lambda)}\\ \overline{g_{2}(\lambda)}\\ \vdots\\ \overline{g_{n}(\lambda)}\end{pmatrix}=\left(\left[\langle f_{i},f_{j}\rangle_{\mu}\right]_{1\leq i,j\leq n}\right)^{-1}\begin{pmatrix}f_{1}(\lambda)\\ f_{2}(\lambda)\\ \vdots\\ f_{n}(\lambda)\end{pmatrix}$.
@@
    The Lemma [18, Lemma 4.2, 4.3] provides formula for obtaining $\langle f_{i},f_{j}\rangle_{\mu}$ for any $1\leq i,j\leq n$ as follows :
    (2.7) $\|f_{i}\|_{\mu}^{2}=\langle f_{i},f_{i}\rangle_{\mu}=c_{i}\zeta_{i}f_{i}^{\prime}(\zeta_{i})$
@@
    (2.8) $\langle f_{i},f_{j}\rangle_{\mu}=\displaystyle\frac{1}{O_{\mu}^{\prime}(\zeta_{i})\overline{O_{\mu}^{\prime}(\zeta_{j})}(1-\zeta_{i}\overline{\zeta_{j}})}.$
@@
Let $\mu$ be a positive Borel measure on unit circle $\mathbb{T}$ defined as $\mu=\delta_{1}+\delta_{\xi}$. The reproducing kernel for $D(\mu)$ is given by:
@@
$+\displaystyle\frac{(z-\xi)(\bar{\lambda}-1)}{\Delta(\xi-1)}+\frac{(z-1)(\bar{\lambda}-\bar{\xi})}{\Delta(\bar{\xi}-1)}\Bigg],~~z,\lambda\in\mathbb{D}$
where $a,b,\Delta\in\mathbb{R}$ are constants and $q(z)$ is a polynomial of degree $2$ whose roots are in $\mathbb{C}\setminus\overline{\mathbb{D}}$.